OpenStudy (anonymous):

find the circumference of a circle whose area is 452.16 square meters. (use 3.146 as the value of pie.)

OpenStudy (anonymous):

area=pi*r^2 circumference=2*pi*r where r is the radius

OpenStudy (bradely):

pir^2=452.16 r^2 =452.16/pi =452.16/3.146 =144 r =12 circumference of a circle 2pir =2pi*12 =24pi =24*3.146 =75.5 m Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Thanks guys! This is really helping me to review for an upcoming test :)